OpenStudy (anonymous):

Solve using matrices. Please show all work :) 3x + 5y = -4 x + 2y = -2

OpenStudy (wolf1728):

I can solve this using Cramer's Rule: http://www.1728.org/cramer.htm 3x + 5y = -4 x + 2y = -2 Let a = 3 b=5 e=-4 c=1 d=2 f=-2 Denominator determinant dn = a•d -c•b = 6 -5 = 1 x determinant = [e•d - f•b] / dn = -4 • 2 - -2•5] / 1 = -8 +10 = 2 y determinant = [a•f -c•e]/1 = [3•-2 -1•-4]/1 = -6 +4 =-2